OpenStudy (anonymous):

At a large school, 110 students were absent on Monday, 100 were absent on Tuesday, and 90 were absent on Wednesday. In total there were 170 students who were absent at least one of these three days. What is the largest number of students who could have been absent for every one of these days?

OpenStudy (anonymous):

Let x be the number of students who were absent for every one of these days. x + m = 110 x + t = 100 x + w = 90 x + m + t + w = 170

OpenStudy (anonymous):

x + t = 100 t = 100 - x x + w = 90 w = 90 - x x + m + t + w = 170 110 + 100 - x + 90 - x = 170 300 - 2x = 170 2x = 130 x = 65

OpenStudy (anonymous):

thx